OpenStudy (anonymous):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10 Answer 2y + 10 -2y - 10 -2y + 10 2y - 10

OpenStudy (anonymous):

-x - 2y = -10 -x = -10 + 2y x = 10 - 2y = -2y + 10 So -2y + 10

jhonyy9 (jhonyy9):

x=10-2y